OpenStudy (anonymous):

solve by completing the square.. PL0X : x²-4x+2=0 otherwise known as x^2-4x+2=0

OpenStudy (lgbasallote):

x^2 - 4x +(4/2)^2 = -2 +(4/2)^2 4/2 ^2 = 4 so... x^2 - 4x + 4 =-2 + 4 x^2 - 4x +4 = 2 (x-2)^2 = 2 x-2 = sqrt 2 x = 2 plus/minus sqrt 2